OpenStudy (anonymous):

Can someone please help me with this one: -1/9(x-27)+1/3(x+3)=x+67

OpenStudy (anonymous):

expand d brackets

OpenStudy (anonymous):

u wud get....2/9x + 4 = x + 67

OpenStudy (anonymous):

equate....put x on 1 side of d = sign an constants on the other therefore it reads, -7/9x = 63

OpenStudy (anonymous):

x = 81